OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line.(7,-1);3x-2y=5

OpenStudy (anonymous):

Since it is parallel, the slope will be the same. and then you ll just need to plug in the values in y = mx + b

OpenStudy (anonymous):

I should make the problem read 3(7)-2(-1)=5 and then solve?

OpenStudy (anonymous):

or maybe it is supposed to be 5=3(7)-2(-1)+b